# Enhanced Sampling for Efficient Learning of Coarse-Grained Machine Learning Potentials

### v1

@@ 4 Conclusion
One limitation of our approach is its dependence on prior knowledge of collective variables (CVs) or reaction coordinates suitable for biasing. In the Müller–Brown and capped alanine benchmarks, the relevant slow modes are well understood, allowing the bias to be applied directly to the coarse-grained degrees of freedom. However, for more complex biomolecular systems, it is challenging to identify such CVs 78, 79, 80. When chosen CVs do not adequately capture the slow dynamics, enhanced sampling may not sufficiently enrich force coverage, limiting the improvement of the resulting models. Machine learning techniques for automated reaction coordinate discovery 81, 82, 83, 84, 85 provide a potential solution, and their integration could facilitate a broader application to high-dimensional CG mappings and larger biomolecules.
Looking ahead, the framework allows for several natural extensions. Biasing could be applied not only along predefined collective variables, but also along arbitrary coarse-grained degrees of freedom, including learned slow coordinates. Additional enhanced sampling methods, such as adaptive biasing force 86, 87, 88, replica exchange or tempering 7, could readily be integrated to further improve efficiency. One could also bias orthogonal degrees of freedom and recover unbiased mean forces with subsequent reweighting strategies 89 (e.g., via Markov state models 90), though the benefit is likely limited, since coarse-grained mappings are typically designed to capture the slowest, most relevant coordinates.
Additionally, an active learning cycle alternating between model training, uncertainty quantification of mean forces, and targeted bias placement would enable systematic sampling of regions with high uncertainty, producing datasets that are both efficient and informative 36, 91, 92. The balance between biased and unbiased simulations can also be optimized, for example, by employing pretraining-finetuning paradigms that take advantage of complementary data sources 93. From a practical perspective, our approach could be used to construct or refine large-scale datasets for training transferable CG MLPs 94, 95, improving transferability across varying thermodynamic conditions and chemical compositions. It could also provide information for generative models that typically lack force supervision 96 or support energy-based neural samplers 97, 59, 98. Overall, we believe that our method represents a fundamental advance over current methodologies and opens new opportunities to tackle outstanding challenges for efficient learning of coarse-gained molecular models.
@@
Details of dataset generation, training procedures, hyperparameters, and additional qualitative results on molecular systems are available in the Supporting Information.

### v2

@@ 4 Conclusion
One limitation of our approach is its dependence on prior knowledge of collective variables (CVs) or reaction coordinates suitable for biasing. In the Müller–Brown and capped alanine benchmarks, the relevant slow modes are well understood, allowing the bias to be applied directly to the coarse-grained degrees of freedom. However, for more complex biomolecular systems, it is challenging to identify such CVs 82, 83, 84. When chosen CVs do not adequately capture the slow dynamics, enhanced sampling may not sufficiently enrich force coverage, limiting the improvement of the resulting models. Machine learning techniques for automated reaction coordinate discovery 85, 86, 87, 88, 89 provide a potential solution, and their integration could facilitate a broader application to high-dimensional CG mappings and larger biomolecules.
Looking ahead, the framework allows for several natural extensions. Biasing could be applied not only along predefined collective variables, but also along arbitrary coarse-grained degrees of freedom, including learned slow coordinates. Additional enhanced sampling methods, such as adaptive biasing force 90, 91, 92, replica exchange or tempering 7, variationally enhanced sampling 93, 94, could readily be integrated to further improve efficiency and transferability. Furthermore, one could relax the requirement of biasing only the mapped coordinates by applying bias to orthogonal degrees of freedom; this would necessitate recovering unbiased mean forces via explicit reweighting strategies 95, 96 (e.g., via Markov state models 97, 98). Such extensions share conceptual roots with established frameworks for transferring potentials between thermodynamic states 99, 100, 35, 101. However, the benefit of biasing orthogonal degrees of freedom is likely limited in this context, since coarse-grained mappings are typically designed to explicitly capture the slowest, most relevant coordinates.
Additionally, an active learning cycle alternating between model training, uncertainty quantification of mean forces, and targeted bias placement would enable systematic sampling of regions with high uncertainty, producing datasets that are both efficient and informative 36, 102, 103. The balance between biased and unbiased simulations can also be optimized, for example, by employing pretraining-finetuning paradigms that take advantage of complementary data sources 104. From a practical perspective, our approach could be used to construct or refine large-scale datasets for training transferable CG MLPs 105, 106, improving transferability across varying thermodynamic conditions and chemical compositions. It could also provide information for generative models that typically lack force supervision 107 or support energy-based neural samplers 108, 61, 109. Overall, we believe that our method represents a fundamental advance over current methodologies and opens new opportunities to tackle outstanding challenges for efficient learning of coarse-gained molecular models.
@@
Details of dataset generation, training procedures, hyperparameters, computational cost, and additional qualitative results on molecular systems are available in the Supporting Information.